OpenStudy (oswaldmurphy):

I just need someone to say correct or not. No explanation unless I am wrong please :D 4x^2 + 6x + 2 = 0 a: 4 b: 6 c: 2

OpenStudy (oswaldmurphy):

x = -6 +- _/6^2 - 4(4)(2)/2(4) x = -6 +- _/36 - 32/8 x = -6 +- _/4/8 x = -6 +- 2/8 +: x = -4/8 x = -1/2 -: x = -8/8 x = -1

OpenStudy (oswaldmurphy):

http://prntscr.com/c21bl8

OpenStudy (oraclethinktank):

You are correct.

OpenStudy (oswaldmurphy):

@Ciarán95 @pooja195 @mathmate

OpenStudy (oswaldmurphy):

So what would be the answer? Its a fill in the blank kinda thing

OpenStudy (oraclethinktank):

x= -1/2, -1

OpenStudy (oswaldmurphy):

It says numerical answers expected..

OpenStudy (mathmate):

@OswaldMurphy Yes, your solution is correct.

OpenStudy (oswaldmurphy):

http://prnt.sc/c21bl8 What would go in the blank @mathmate

OpenStudy (oraclethinktank):

Wouldn't it be x= -1/2, -1?

OpenStudy (oswaldmurphy):

Numerical Answers.

OpenStudy (mathmate):

Typically when a quadratic can be factored, we mean it can be factored using rational numbers. The next question to ask if a quadratic can be factored is therefore "is the solution rational?". @OswaldMurphy which part of the quadratic formula could introduce irrational numbers?

OpenStudy (oswaldmurphy):

Squaring

OpenStudy (mathmate):

rational number means a number that can be represented by a fraction, such as 22/7, -4, etc. Irriational numbers canNOT be represented by a fraction, such as \(\pi,\sqrt{3}\), etc. Do you think the quadratic formula below can introduce an irrational number, and where? \(\Large \frac{-b\pm\sqrt{b^2-4ac}}{2a}\)

OpenStudy (oswaldmurphy):

The division

OpenStudy (mathmate):

Please show me an example of a division that makes irrational numbers.

OpenStudy (mathmate):

Recall that 22/7 is a rational number.

OpenStudy (oswaldmurphy):

Yes.